January 25, 2017 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, January 25, 2017
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates
  3. Data Validation
  4. GlueX Detector Performance Document

Upcoming Calibrations

  • Add timing information to FCAL clustering/inner ring calibrations? (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • TAGH alignment (Nathan)
  • Tagger energy map (Richard J.?)
  • TOF alignment (Beni/Simon)


  1. Announcements
  2. Collaboration Meeting
  3. Calibration Updates
  4. Simulations
    • sim1.2 progress
  5. AOB


Attending: Sean (NU); Simon, Mark D., Adesh, Eugene, Alex A., Thomas, Brad, Nathan, Mark I. (JLab); Curtis, Mike S., Will M., Naomi (CMU); Matt S. (IU); Justin (W&M); Mahmoud (FIU)

  1. Announcements
    • Beam is expected next week.
  2. Collaboration Meeting
    • Some discussion of talks for Saturday session: Sean and Mark I. will give calibration/offline overview talks, detector talks should cover performance & simulation plans, Mike S. will give alignment talk in tracking session.
    • Other suggestions welcome, Sean will look through his notes.
  3. Calibration Updates
    • Sean presented a brief update on the timing calibrations from the last run. The TOF calibrations look good, with resolutions approaching their values from last Spring. All the others have values from the calibration before the TOF recalibration, and the per-channel timing calibrations are not good.
      • Studies of the SC have shown a couple curious features: (1) the propogation time corrections have clearly changed, in some cases by large amounts; (2) some paddles seem to have been calibrated to the wrong RF bin and have a +-2 ns offset. For (2) it is not clear why this is the case, there may be some theta-dependent effect. For (1) the cause is even more unclear, since the flash algorithm was supposedly essentially unchanged. This is a sign that we carefully need to look at the calibrations for all f250-dependent detectors (i.e. all but the drift chambers).
      • Most detectors are out of alignment with the RF in ~+-1 ns range.
      • Sean will average out the SC offsets and rerun calibrations to hopefully get some better alignment. Some more BCAL calibrations will be run, in collaboration with Mark Dalton.
      • Curtis asked what beam frequency we will get in the upcoming run, and Eugene said that it would be at 250 Mhz since the RF separator will be used. Note that this last run was at 500 Mhz.
      • Matt pointed out that there is a difference in how the pedestals were calculated between the firmware, which is small but could be meaningful.
    • The Calibrations train is being updated for the new run, developments delayed due to resolving calibration issues.
    • FCAL - Adesh has been working on the pi0 calibration. He started with gains of 1 and the same HV settings as the spring 2016 ended with. At the start of the run, he noticed that the pi0 peak shifted from 135 MeV to 130 MeV, the reason is not understood. After the 7th gain iteration, the pi0 width has reduced from 8.7 MeV to 7.5 MeV, compared to 6.7 MeV in the spring 2016 run. New calibrations have been added to the CCDB.
      • Eugene pointed out that there could be some difference in the pi0 energy spectrum due to various conditions, e.g. changing trigger thresholds. Adesh will look at this.
      • Sean asked about changes in pileup or other backgrounds, Adesh hadn't seen anything
      • Mike S. suggested to run the emulation code for the old and new firmware over the data taken in the fall and compare the results.
      • Adesh had mentioned to Sean that he doesn't have enough data to do a precision timing calibration of the FCAL. Approximately several days of steady production data would be needed.
    • BCAL - Will has been working on the pi0 calibrations. The integration window was changed from the spring to the fall run, and with an eyeballed change to the gain factors, the pi0 peak stood at 140 MeV at the beginning of the run. Recalibration has brought the peak back to 135 MeV, but the resolution for a >1 GeV photon energy cut is 16 MeV, compared to 7.5 MeV in the previous run.
      • Mark Dalton: We are going through and looking at all of the other calibrations. The attenuation length and gain ratios are on my list. The effective velocities should be looked at as well.
      • Integration window settings will likely be discussed at tomorrow's calorimetry meeting. The safe choice is to go back to the spring 2016 integration window. Justin has been doing some studies on correcting for saturation in the BCAL with a method that works better for longer windows. The BCAL is sensitive to pedestal fluctuations, though.
      • Matt asked about the possibility of some artificial timing jitter, or other unusual effect, being introduced somewhere. He also suggested running some tests using LED data. Mark pointed out that the firmware studies Sean has done seem to suggest the firmware is doing what it is supposed to, and that he had done such studies last spring with the old firmware in order to determine the new integration windows, but it might be worth doing them again with the new firmware as a check.
    • CDC - Naomi has been performing gain calibrations with the latest data. She found a good one with low pressure (high gain) to do an individual straw calibration, and was able to calibrate the overall scale for runs with a HV of >= 2125 V. The operating voltage for the next run is being studied.
    • Alignment - Mike's efforts to implement alignment with Millipede have been going forward. He is currently generating some simulations to understand what level of improvements can be expected with straight tracks.
    • SC - Status was discussed in previous timing discussion, nothing new, Mahmoud will look at propogation times after timing recalibration.
    • TOF - New HV settings for each channel have been obtained. Recalibration of the TOF is the first priority for the upcoming run. These settings should recover the efficiency loss in the full paddles near the beam hole noted by Sean in yesterday's TOF meeting.
    • TAGM - Alex B. not at meeting, in the RC meeting he mentioned that he hasn't been able to analyze data taken to check the newly installed fibers. Mark I. had a similar problem the with SC, this is being resolved in discussion with experts. Sean will follow up with Alex about a calibration plan for the TAGM that he had requested.
    • TAGH - Nathan has been working on the ADC/TDC timing offsets. ~20% of channels have been misaligned by 1-2 2ns beam buckets. This is an improvement over the previous run, where much larger shifts were seen, but some improvement in the calibration procedure may be able to be made.
  4. Simulations
    • At the last offline meeting, it was noted that only a few percent of the sim1.2 jobs had actually finished. Most of the time was spent simulating EM background, which is not important for the acceptance and other studies that these simulations were needed for. The decision to restart production without EM background was made.
    • Sean mentioned that there are some small improvements to mcsmear that have been put on the shelf to deal with other problems.
    • Implementation of mix-in backgrounds to simulation is a major goal for this spring.
    • Mark I. reported that the OSG-gateway node has been ordered, does not exist at the lab yet.
  5. Matt mentioned that the FCAL LED skim files are empty. Sean suggested using tonight's cosmic run to see why the offline skimming is missing these and other front-panel triggers.